OpenStudy (anonymous):

Area of a rectangle is 336 m^2. Ratio of Width and Length is 3:7. Find Perimeter

OpenStudy (anonymous):

So I know Area = W x L and that Perimeter = 2 ( W x L ) ...

OpenStudy (anonymous):

I mean the perimeter = 2 ( W + L )

OpenStudy (nikato):

Since the ratio is 3:7 lets say. 3x for width and 7x for length So area is 3x(7x)=336 Can you solve for x?

OpenStudy (nikato):

@Gulivr You'll get 21x^2=336 21x^2 -336=0 Factor a 21 21(x^2 -16)=0 So x^2-16=0 And you can factor the left side (x-4)(x+4)=0 So x=4 or -4 But since lengths can't be negative x=4

OpenStudy (nikato):

And if we go back to the beginning, We said width=3x. And length=7x So plug in x Width=12. And length =28 And now you should be able to find the perimeter